# Efficient Transformer-based Hyper-parameter Optimization for Resource-constrained IoT Environments

### v1

@@ I Introduction
Convolutional Neural Networks (CNNs) are the staple implementation of neural networks (NNs) for image classification and object detection. The progress in this field is attributed to the increased complexity of CNN architectures in terms of the type of connections among different layers and their depth, which increased the computational demands. As a result, the accuracy of these models depends on the number and the type of layers, the connections between these layers, and the parameters assigned to each layer [1]. In addition to accuracy, other important factors that dictate the design of CNNs include training time, inference time, and energy consumption, which are emphasized in the resource-constrained Internet of Things (IoT) environment. The number of possible choices for designing CNNs is extremely large, which is prohibitive for manual search. These circumstances have promoted the automated CNN architecture search, achieved through the field of Neural Architecture Search (NAS).
@@
The NAS field has recently seen great progress, due to the incorporation of Reinforcement Learning (RL) agents to search for the best CNN configurations. The RL’s appeal stems from its generalizability of hyper-parameter (HP) combinations via function approximation and the trial-and-error approach, which has the potential to reduce the computational demand of NAS [2]. This field is dominated by the literature that either builds CNNs from scratch using basic NN layers or a pre-defined collection of these layers, or optimizes the HPs of an already-existing CNN.
@@
Autonomous vehicles (AV) are part of the envisioned IoT applications that utilize edge servers of limited computational resources. The realization of AVs heavily depends on image classification models deployed on these servers. As such, the disparate computational capabilities of edge servers are prohibitive for the prolonged execution of large models. Model partitioning is key to resolving these limitations toward fulfilling the promise of AVs [3]. This requires gathering insights into the contribution of each CNN model’s layers to the classification results. In current implementations, this vision is stalled by multiple oversights. The sequential nature of model generation results from the dependence between layer combinations. This dependence can be reflected by the recurrent structures of Recurrent Neural Networks (RNNs). Therefore, RNN-based controllers constructing CNN models dominated the RL-based NAS field [4, 5]. The one-at-a-time processing of inputs inhibits their parallelization, resulting in their insurmountable computational footprint that is ill-fitted for the resource-scarce vehicular environment. On a different note, these controllers are built to receive a reward when a terminal condition is encountered, masking the contribution of each layer to prediction results. Within the constraints of computational resources and model partitioning requirements, the state-of-the-art (SOTA) approaches are limited by their impracticability and lack of transparency, hindering their deployment in real-world environments.
@@
This paper presents Transformer-based Reinforcement Learning Hyper-parameter Optimization (TRL-HPO) that alleviates the shortcomings of the current methods. The TRL-HPO addresses the transparency and long-computational times of RL-based solutions with competitive performances. The Transformer architecture overshadows the RNN-based methods by integrating the multi-headed self-attention mechanism [6]. The TRL-HPO controller equipped with the attention mechanism allows us to gain insights from the layer-generation process, enhancing its interpretability. Moreover, the transformers’ architecture supports parallelization, which addresses the long computation times of its RNN-based counterparts [6]. The reward is produced with every generated layer to showcase improvement instead of waiting for termination conditions to unfold, a condition matching CNN partitioning requirements, as depicted in Figure 1. The contributions of this paper are as follows:
@@
  • Propose a novel Hyper-parameter Optimization (HPO) process named TRL-HPO that is the first to combine transformer architecture and actor-critic RL;
@@ II Related Work
Bayesian Optimization (BO), Evolutionary Search (ES) algorithms, and RL agents are the three main tools for HPO implementation. BO and ES methods are limited by their assumptions [1] and lack of generality [7], which favors the application of RL techniques. The works of Baker et al. [5] and Zoph et al. [4] are the first to propose incorporating RL methods into generating CNN models. The former work utilized RNN-based controllers using an off-policy RL algorithm to sample CNN architectures while the latter utilized a value-based Q-learning approach. The results of these seminal works highlighted the trade-off between the running time of HPO methods and the accuracy of the obtained models.
@@
The shortcomings exhibited by the integration of RL into the HPO process in these implementations hinder their deployment in IoT environments. First, the adopted models lack the transparency that shows the interdependence between different layers. Second, some of these approaches take considerable time to converge, which is a privilege that is not always available in resource-constrained environments. Third, some works focus on optimizing the HPs of a specific layer, reducing the state space at the expense of the layer’s diversity. Lastly, many of the works include prior knowledge instilled into the layer-generation process, such as the addition of dropout layers [4, 5]. These limitations are addressed in this paper by proposing a transformer-based RL controller and the reward function formulation. The multi-headed attention mechanism facilitates the training process and adds transparency to the model generation process by analyzing the attention values. On the other hand, the reward function reflects the contribution of each layer to enhancing the classification results, a condition favorable to CNN model partitioning.
[…]
[scale=0.45]architecture_v3.PNG
@@ III Methodology: TRL-HPO
This section explains the building blocks of TRL-HPO, including the transformer and RL’s actor-critic architectures, which are shown in Figure 1, and the motivation for this combination toward a more efficient and transparent HPO process.
@@ III-A Transformer
The vanilla transformer is a sequence-to-sequence model, which avoids the recurrence structures of RNNs by integrating the innovative self-attention mechanism [6]. This architecture enables parallelization accelerating the training of transformer models. The building blocks of transformers include attention mechanism, multi-headed attention, position-wise feed-forward network, and positional encoding (PE). The transformers follow the encoder-decoder structure thereby each of these structures is formed by stacking combinations of these identical layers. The ordering of these layers with respect to the encoder and decoder is depicted in Figure 1.
The Positional Encoding computes the sequence’s order and is added to the input embedding of the encoder and decoder stacks so that the order is incorporated into the input and output data. To obtain a unique order, a sinusoidal function takes as inputs the position and the embedding dimensions. This structure extrapolates to sequences that are longer than the ones encountered during training [6]. The self-attention module calculates the weights assigned by each input sequence to other sequences. This way the recurrence structure is avoided because self-attention reflects the dependencies between input and specific output and improves the modelling of long-range dependencies [6]. As such, the attention mechanism resembles a fully connected layer (FCL), such that the weights reflect the pairwise relationship from previous inputs. To fully exploit this method, the transformer includes multi-headed attention that replicates the attention calculation. This process generates different representations of the input data at different positions. The calculation of these attention matrices is independent and facilitates their parallelization. Lastly, a feed-forward neural network is used as the subsequent layers to the multi-head attention mechanism.
@@ III-B Actor-Critic Reinforcement Learning
RL teaches an agent to perform a task by accumulating experiences from interacting with its environment. Its actions are refined based on a reward function that signals the utility of these actions [12]. In connection to the HPO process, the RL’s experimentation with the absence of any prior knowledge mirrors the inability to hypothesize the best combination of basic CNN layers. The only knowledge for the HPO process relates to the order of the stacked layer that starts with grid-like inputs and ends with a fully connected layer (FCL). As such, the combination of unknown orientation and trial-and-error experimentation matches the requirements of the HPO process.
The RL methods can be split into value-based and policy-based methods. The value-based method estimates the quality of a state-action pair using a value function and optimizes this value iteratively using actions that maximize it. The high variability of value estimations and their under-performance in scenarios with continuous action spaces promotes policy-based methods [12]. The policy-based methods focus on optimizing the policy to maximize the expected cumulative rewards [12]. However, the latter methods suffer from sample inefficiency, which is an asset of value-based methods. To complement these methods’ advantages, an actor-critic approach combines both policy-based (actor) and value-based (critic) methods [12]. Projected to the HPO problem, the actor outputs an action that maps to a layer and its HPs. The critic evaluates this action by outputting a value showing the quality of the taken action [12].
@@ III-C Transformers and Actor-critic RL
As the name suggests, the proposed framework, Transformer-based Reinforcement Learning HPO (TRL-HPO) is the convergence of transformers and actor-critic RL. The actor and the critic are implemented using the transformer architecture. The steps involved in the actions’ generation, CNN model construction, and their evaluation are depicted in Figure 1. The combination of TRL-HPO enables the RL agent to harness the strengths of transformers to the benefit of the HPO process. This new framework is the first trial to integrate transformers into the HPO process, opening a new frontier toward the exploitation of this novel architecture to address a lingering problem in the field of ML. This experiment does not plan to only exhibit the viability of transformers’ implementation to HPO, but it rather analyzes the suitability of this architecture to the HPO process.
We analyze TRL-HPO based on its inherent advantages and its benefits versus the SOTA approaches. In terms of the framework, two main advantages are reaped. The first is that TRL-HPO builds models from scratch, rendering it a general-purpose method. The proof of concept is constrained to a single use case; however, the framework can be expanded to any problem that utilizes DNN models. The second concerns the definition of the reward function that is progressively updated with each stacked layer. While this definition incurs an extra computational footprint by training progressively deeper models, it facilitates understanding the contribution of each layer. This feature fits the requirements of environments that seek to balance resource constraints and accuracy objectives. The transformer architecture results in two inherent advantages to the HPO process. The first is connected to the parallelization enabled by the multi-headed attention mechanism, which reduces the running time of the HPO process. The second benefit stems from the attention mechanism that showcases the relationships between generated layers. A by-product of this feature is the inference of the combination of layers that produce better results, which enhances the transparency of the HPO process.
@@ IV Implementation Details
The evaluation of TRL-HPO is conducted on the MNIST dataset [13]. This dataset is a large database of handwritten digits, from 0 to 9, containing 60,000 training images and 10,000 testing images. Each image is a greyscale image of 28 $\times$ 28 dimensions. This dataset serves as a good proof-of-concept for the proposed framework, due to its availability and limited complexity. The former factor enables a fair comparison with other SOTA approaches, such as [5, 8]. The latter factor facilitates training CNN models in a limited time; an advantage when working with hardware constraints. The comparison with SOTA approaches is based on the convergence time to the best solution (hrs), the classification accuracy (%), and the best CNN classification model (%) obtained by each SOTA approach upon the completion of TRL-HPO process (AccTime). We compare TRL-HPO with Baker et al. [5] and Neary [8], which are the only RL-based solutions that build CNN models from scratch.
@@
The generated CNN models are constrained to 6 layers, formed as a combination of the three basic layers: CNN layer, FCL layer, and Maximum Pooling (MaxPool) layer. The set of HPs used in the experimental procedure is summarized in table I. Stacking a combination of these layers adapts to the input structure. In connection to the RL agent, the design of the state and action space and the reward function is imperative to drive the RL agents’ action generation, which requires defining the transformers’ input sequence.
@@
The reward function represents a layer’s contribution towards improving the classification results, which means that the DNN model’s first layer produces the highest rewards compared to any subsequent layers. Each layer is represented as a combination of the representation of the layer itself and its HP, and the performance of the obtained model on the MNIST’s validation set. The generated layers’ representation is obtained from the output layer of the RL’s actor, summarized in four values. These values represent the action space that is mapped first to a layer and then to that layer’s HPs. The model’s performance is represented with 32 values, each calculating the model’s accuracy results on each validation set’s batch of size 16. The heterogeneity in the dimensions of the action space and the performance requires a mapping to a uniform representation. This goal is achieved using a static NN that takes these inputs and produces a uniform output of 64 values, referred to as Intermediate Model Representation (IMR). When a layer and its performance are obtained, the state space, reflecting the current state of the environment, should also change. Since the state space represents a sequence of layers, the index corresponding to the generated layer is updated with the IMR. With more layers, each index is updated with a new representation. This way the state space reflects the two important pieces of information for each layer, its HPs, and its performance. This process was touched upon in Figure 1.
@@
The stopping criteria for the RL agent involve either generating more than 6 layers, minimal improvement in the performance with the addition of layers (0.001), or accuracy below 60%. These criteria are defined to avoid any unnecessary generation of models. A Deep Deterministic Policy Gradient (DDPG) [14] handles the continuous action spaces in this environment. The critic and the actor consist of target and online subnets to avoid radical changes in critic and actor updates [14]. The training of RL agents involves exploration and exploitation of the gained knowledge. In DDPG, the agent explores by adding random noise to the sampled actions, so that the agent can experiment with different combinations of layers and obtain their performance results. To remove data correlations for the RL agent inherent in the sequential structure of the data, experience replay (ER) [14] buffer is utilized to store data during the exploration stage.
@@
The training process is realized over several episodes. In each episode, 10 full models (i.e., until a stopping criterion is met) are generated via the utilization of CPU parallelization. Once the ER buffer is full, five RL optimization rounds are implemented in each episode. The number of episodes and the size of the ER buffer are important parameters for the HPO process. Their optimization is imperative to obtain models with good performances in a short period. During the HPO process, the training set consists of 20,000 images while the validation set includes 10,000 images out of the original 60,000 images. The experimental procedure is conducted on SHARCNET’s Graham cluster on a node with one V100 GPU, 64 Gbs of RAM, and 12 cores, such that each experiment does not exceed 7 days to avoid long queuing times. The implementation will be available on a GitHub repository upon publication.

### v2

@@ I Introduction
Convolutional Neural Networks (CNNs) are the staple implementation of neural networks (NNs) for image classification and object detection. The progress in this field is attributed to the increased complexity of CNN architectures in terms of the type of connections among different layers and their depth, which increased the computational demands [1]. Therefore, the accuracy of these models depends on the number and the type of layers, the connections between these layers, and the parameters assigned to each layer [2]. Other important factors dictating the design of CNNs include training time, inference time, and energy consumption, emphasized in the resource-constrained Internet of Things (IoT) environment. The number of possible choices for designing CNNs is extremely large, which promoted the automated CNN architecture search, achieved through Neural Architecture Search (NAS).
@@
The NAS field has recently seen great progress, due to the incorporation of Reinforcement Learning (RL) agents to search for the best CNN configurations. The RL’s appeal stems from its generalizability of hyper-parameter (HP) combinations via function approximation and the trial-and-error approach, which can reduce the computational demand of NAS [3]. This field is dominated by the literature that either builds CNNs from scratch using basic NN layers or a pre-defined collection of these layers or optimizes the HPs of an already-existing CNN.
@@
Autonomous vehicles (AV) are part of the envisioned IoT applications that utilize edge servers of limited computational resources. The realization of AVs heavily depends on image classification models deployed on these servers. Their disparate computational capabilities are prohibitive for the prolonged execution of large models. Model partitioning is key to resolving these limitations toward fulfilling the promise of AVs [4]. This requires gathering insights into the contribution of each CNN model’s layers to the classification results. In current implementations, this vision is stalled by multiple oversights. The sequential nature of model generation results from the dependence between layer combinations, which can be reflected by the recurrent structures of Recurrent Neural Networks (RNNs). Therefore, RNN-based controllers constructing CNN models dominated the RL-based NAS field [5, 6]. The one-at-a-time processing of inputs inhibits their parallelization, resulting in their insurmountable computational footprint that is ill-fitted for the resource-scarce vehicular environment. On a different note, these controllers are built to receive a reward when a terminal condition is encountered, masking the contribution of each layer to prediction results. Within the constraints of computational resources and model partitioning requirements, the state-of-the-art (SOTA) approaches are limited by their impracticability and lack of transparency, hindering their deployment in real-world environments.
@@
This paper presents a Transformer-based Reinforcement Learning Hyper-parameter Optimization (TRL-HPO) model that alleviates the shortcomings of the current methods. The TRL-HPO addresses the transparency and long-computational times of RL-based solutions with competitive performances. The Transformer architecture overshadows the RNN-based methods by integrating the multi-headed self-attention (MHSA) mechanism that enables parallelization [7], addressing the long computational times. The TRL-HPO controller equipped with the attention mechanism allows us to gain insights from the layer-generation process, enhancing its interpretability. The reward is produced with every generated layer to showcase improvement instead of waiting for termination conditions to unfold, a condition matching CNN partitioning requirements, as depicted in Figure 1. The contributions of this paper are as follows:
@@
  • Propose a novel Hyper-parameter Optimization (HPO) process named TRL-HPO that is the first to combine transformer architecture and actor-critic (AC) RL;
@@ II Related Work
Bayesian Optimization (BO), Evolutionary Search (ES) algorithms, and RL agents are the three main tools for HPO implementation. BO and ES methods are limited by their assumptions [2] and lack of generality [8], which favors RL techniques. The works of Baker et al. [6] and Zoph et al. [5] are the first to propose incorporating RL methods into generating CNN models. The former work utilized RNN-based controllers using an off-policy RL algorithm to sample CNN architectures while the latter utilized a value-based Q-learning approach. The results of these seminal works highlighted the trade-off between the running time of HPO methods and the accuracy of the obtained models.
@@
The shortcomings of RL integration into the HPO process hinder their deployment in IoT environments. First, the adopted models lack the transparency that shows the interdependence between different layers. Second, the long convergence times are prohibitive for deploying these models in resource-constrained environments. Third, some works focus on optimizing the HPs of a specific layer, reducing the state space at the expense of the layer’s diversity. Lastly, many works include prior knowledge in the layer-generation process, such as the addition of dropout layers [5, 6]. This paper addresses these limitations by proposing a transformer-based RL controller and the reward function formulation. The MHSA facilitates the training process and adds transparency to the model generation by analyzing the attention values. On the other hand, the reward function reflects the contribution of each layer to enhancing the classification results, which favors CNN model partitioning.
@@ III Methodology: TRL-HPO
This section explains the building blocks of TRL-HPO, including the transformer and RL’s AC architectures, which are shown in Figure 1, and the motivation for this combination toward a more efficient and transparent HPO process.
@@ III-A Transformer
The vanilla transformer is a sequence-to-sequence model, which avoids the recurrence structures of RNNs by integrating the innovative self-attention mechanism [7]. This architecture enables parallelization accelerating the training of transformer models. The building blocks of transformers include attention mechanism, multi-headed attention (MHSA), position-wise feed-forward network, and positional encoding (PE). The transformers follow the encoder-decoder structure thereby each of these structures is formed by stacking combinations of these identical layers, as depicted in Figure 1.
The PE computes the sequence’s order and is added to the input embedding of the encoder and decoder stacks so that the order is incorporated into the input and output data. To obtain a unique order, a sinusoidal function inputs the position and the embedding dimensions. The self-attention module calculates the weights each input sequence assigns to other sequences. This way the self-attention reflects the dependencies between input and specific output and improves the modelling of long-range dependencies [7]. As such, the attention mechanism resembles a fully connected layer (FCL), whereby the weights reflect the pairwise relationship from previous inputs. To fully exploit this method, the transformer includes MHSA that splits the attention calculation among different heads of the embedding that can be calculated independently, facilitating its parallelization. Lastly, a feed-forward neural network (FFNN) is the subsequent layer to the MHSA.
@@ III-B Actor-Critic Reinforcement Learning
RL teaches an agent to perform a task by accumulating experiences from interacting with its environment. Its actions are refined based on a reward function that signals the utility of these actions [13]. The RL’s experimentation with the absence of prior knowledge mirrors the ignorance of the best combination of basic CNN layers. The only knowledge for the HPO process relates to the order of the stacked layer that starts with grid-like inputs and ends with an FCL. The combination of unknown orientation and trial-and-error experimentation matches the requirements of the HPO process.
The RL methods can be split into value-based and policy-based methods. The value-based method estimates the quality of a state-action pair using a value function and optimizes this value iteratively using actions that maximize it. The high variability of value estimations and their under-performance in scenarios with continuous action spaces promotes policy-based methods [13]. The policy-based methods optimize the policy to maximize the expected cumulative rewards [13]. However, the latter methods suffer from sample inefficiency, which is an asset of value-based methods. To complement these methods’ advantages, an actor-critic (AC) approach combines both policy-based (actor) and value-based (critic) methods [13]. Projected to the HPO problem, the actor outputs an action that maps to a layer and its HPs. The critic evaluates this action by outputting a value showing the action’s quality [13].
@@ III-C Transformers and Actor-critic RL
The proposed framework, Transformer-based Reinforcement Learning HPO (TRL-HPO) is the convergence of transformers and AC RL. The actor and the critic are implemented using the transformer’s decoder architecture. The steps involved in the actions’ generation, CNN model construction, and their evaluation are depicted in Figure 1. The combination of TRL-HPO enables the RL agent to harness the strengths of transformers to benefit the HPO process. This new framework is the first trial to integrate transformers into the HPO process, opening a new frontier toward the exploitation of this novel architecture to address a lingering problem in the field of ML. This experiment evaluates the viability and suitability of this architecture to the HPO process.
We analyze TRL-HPO based on its inherent advantages and benefits versus the SOTA approaches. Two main advantages are reaped. The first is that TRL-HPO builds models from scratch, rendering it a general-purpose method. The proof of concept is constrained to a single use case; however, the framework can be expanded to any DNN-related problem. The second concerns the definition of the reward function that is progressively updated with each stacked layer. While this definition incurs an extra computational footprint by training progressively deeper models, it facilitates understanding the contribution of each layer. This feature fits the requirements of environments that seek to balance resource constraints and accuracy objectives.
The transformer architecture results in two main advantages to the HPO process. The first relates to handling long sequences, which benefits the transferability of TRL-HPO from small to larger datasets. The second benefit relates to the MHSA, which facilitates the parallelization of TRL-HPO and reduces its running time. Additionally, MHSA reveals the relationships between generated layers, which allows the inference of the combination of layers that produce better results, enhancing the transparency of the HPO process.
@@ IV Implementation Details
The evaluation of TRL-HPO is conducted on the MNIST dataset [14], which is a large database of handwritten digits, from 0 to 9, containing 60,000 training and 10,000 testing images. Each image is a greyscale image of 28 $\times$ 28 dimensions. The availability and limited complexity of MNIST can serve as a good proof-of-concept for TRL-HPO. The former factor enables a fair comparison with other SOTA approaches, such as [6, 9]. The latter factor facilitates training CNN models in a limited time; an advantage when working with hardware constraints. The comparison with SOTA approaches is based on the convergence time to the best solution (hrs), the classification accuracy (%), and the best CNN classification model (%) obtained by each SOTA approach upon the completion of TRL-HPO process (AccTime).
Reproducibility is a major issue that plagues the HPO process [2]. With this factor in mind, the choice of competing methods depended on two conditions: (1) The availability of source code that enables a fair comparison with TRL-HPO and dispels any introduced biases by implementing a method from scratch. (2) The similarity in the experimental procedure for building CNNs from scratch using an RL-based method and the availability of results applied to the MNIST dataset. Therefore, we compare TRL-HPO with Baker et al. [6] and Neary [9], which abide by at least one of these conditions.
@@
The generated CNN models are constrained to 6 layers, formed as a combination of the three basic layers: CNN layer, FCL layer, and Maximum Pooling (MaxPool) layer. The set of HPs used in the experimental procedure is summarized in table I. Stacking a combination of these layers adapts to the input structure. The state and action space design and the reward function are imperative to drive the RL agents’ action generation, which requires defining the transformers’ input sequence.
@@
The reward function represents a layer’s contribution towards improving the classification results, which means that the DNN model’s first layer produces the highest rewards compared to any subsequent layers. Each layer is represented as a combination of the representation of the layer itself and its HP, and the performance of the obtained model on the MNIST’s validation set. The generated layers’ representation is obtained from the output layer of the RL’s actor, summarized in four values. These values represent the action space mapped first to a layer and then to that layer’s HPs. The model’s performance is represented with 32 values, each calculating the model’s accuracy results on each validation set’s batch of size 16. The heterogeneity in the dimensions of the action space and the performance requires a mapping to a uniform representation. This goal is achieved using a static NN that takes these inputs and produces a uniform output of 64 values, referred to as Intermediate Model Representation (IMR). When a layer and its performance are obtained, the state space, reflecting the current state of the environment, should also change. Since the state space represents a sequence of layers, the index corresponding to the generated layer is updated with the IMR. With more layers, each index is updated with a new representation. This way the state space reflects the two important pieces of information for each layer, its HPs, and its performance. This process was touched upon in Figure 1.
@@
The stopping criteria for the RL agent involve generating more than 6 layers, minimal improvement in the performance with the addition of layers (0.001), or accuracy below 60%. These criteria are defined to avoid any unnecessary generation of models. A Deep Deterministic Policy Gradient (DDPG) [13] handles the continuous action spaces in this environment. The critic and the actor consist of target and online subnets to avoid radical changes in critic and actor updates [13]. In DDPG, the agent explores by adding random noise to the sampled actions, so that the agent can experiment with different combinations of layers and obtain their performance results. To remove data correlations for the RL agent inherent in the sequential structure of the data, experience replay (ER) [13] buffer is utilized to store data during the exploration stage.
The actor and the critic follow the transformer architecture, depicted in Figure 1. Two main differences highlight these components’ distinctive roles: the transformer’s output and the learning rate (lr). The actor’s output is in the [0, 1] range mapped to a layer, while the critic’s is in the [-1, 1] range representing the state’s Q-value. Regarding the lr, it is recommended that the actors’ (1e-5) be slower than the critics’ (1e-4) [13]. The input embedding’s dimension for each input space is equal to 64. The embeddings are inputs to multiple encoding layers, equivalent to 2 in our implementation, each constituting a Transformer block, represented by the MHSA and FFNN. Each of these blocks has the same number of input and output dimensions. The definition of MHSA and FFNN requires highlighting two values, number of heads equal to 4 and expansion factor equal to 4.
@@
The training process is realized over several episodes. In each episode, 10 full models are generated via CPU parallelization. Once the ER buffer is full, five RL optimization rounds are implemented in each episode. The number of episodes and the size of the ER buffer are important parameters for the HPO process. Their optimization is imperative to obtain models with good performances in a short time. During the HPO process, the training set consists of 20,000 images while the validation set includes 10,000 images out of the original 60,000 images. The experimental procedure is conducted on SHARCNET’s Graham cluster on a node with one V100 GPU, 64 Gbs of RAM, and 12 cores, such that each experiment does not exceed 7 days to avoid long queuing times. The implementation is available on the GitHub repository [^1] https://github.com/Western-OC2-Lab/TRL-HPO.